Airplane A travels 1000 mi at a certain speed. Airplaine B travels the same distance in 1 hr less time but travels 50 mph faster. Find the time that Airplane B travels.
Answer by mananth(16949)

Airplane Time = 1000/x where x is the speed of the plane
Airplane B Time = 1000/x+50 ( 50 mph faster)
..
Time taken by airplane A - Time by Airplane B =1
1000/x - 1000/ x+50=1
LCM = x(x+50)
1000(x+50)- 1000x= x(x+50)
1000x+50,000 - 1000x=x^2 +50x
x^2+50x-50000=0
x^2+250x-200x-50000=0
x(x+250)-200(x+2 b50)=0
(x+250)(x-200)=0
x= 200 mph speed of plane A
speed of plane B = 250 mph
distance traveled = 1000miles
time it traveled = 1000/250 = 4 hours